I updated the SRM payload control.
* I found that the IMP damping loops is not stable. I updated IM_DAMP_P
* IMV seems to have a coupling between pitch. It slowly oscillate 4.7 Hz motion. So I implemented the notch in IM_DAMP_V.
It might be because of the one of the OSEM that is at the edge of the linear range.
It should be solved by diagnalization eventually. This is just a temporary solution.
* OL hierarchical controls are updated in the similar way as other SRs.
* I also implemented the TML control, which has gain 0.03-3 Hz, since I found the report that sometimes L is also noisy.
Memo for further update:
* There are peakings at around 0.2 Hz, where is the cross-over frequency of TM loop and IM loop. Can be reduced by the filtertuning. (Easiest solution is just reducing the IM OLcontrol gain a bit.)
* Also, the cross-over frequencies could be lower so that they are off from the micro-seismic region.
I have thought the cross-over frequencies should be as high as possible (i.e., control the optic via IM as much as possible).
However, it turns out to be difficult to make the cross-over frequneyc higher than about 0.2 Hz due to some peaks. In this case, I am now thinking that the cross-over frequency can be lower than 0.1 Hz.
Anyway, the loops look to reduce the RMS by factor of 2. At some frequencies, it has the gain more than 10.